Calculate in ev the energy of a quantum of light wavelength 53000A
1
Expert's answer
2019-04-25T09:22:19-0400

We can find the energy of a quantum of light from the formula:


"E = \dfrac{hc}{\lambda},"

here, "E" is the energy of a quantum of light, "h = 4.135 \cdot 10^{-15} eV \cdot s" is Plank's constant, "c = 3 \cdot 10^8 m/s" is the speed of light, "\lambda = 53000 \text{\AA}" is the wavelength of the light.

Then, we get:


"E = \dfrac{hc}{\lambda} = \dfrac{4.135 \cdot 10^{-15} eV \cdot s \cdot 3 \cdot 10^8 \dfrac{m}{s}}{53000 \text{\AA} \cdot \dfrac{1 m}{10^{10} \text{\AA} }} = 0.234 eV."

Answer:

"E = 0.234 eV."